# literature.bib

### v1

@@ Abstract
In this paper, we analyze a recently proposed algorithm for the problem of sampling from probability distributions $\mu^{\ast}$ on $\mathbb{R}^{d}$ with Lebesgue density of the form $\mu^{\ast}(x)\propto\exp(-f(Kx)-g(x))$, where $K$ is a linear operator and $f,g$ convex and non-smooth. The algorithm is a generalization of the primal-dual hybrid gradient (PDHG) convex optimization algorithm to a sampling scheme. We analyze the method’s continuous-time limit, an SDE in the joint primal-dual variable. We give mild conditions under which the corresponding Fokker-Planck equation converges to a unique stationary state, which however does not concentrate in the dual variable and consequently does not have $\mu^{\ast}$ as its primal marginal. Under a smoothness assumption on $f$, we show that the scheme converges to the purely primal overdamped Langevin diffusion in the limit of small primal and large dual step sizes. We further prove that the target can never be the primal marginal of the invariant solution for any modification of the SDE with space-homogeneous diffusion coefficient. A correction with inhomogeneous diffusion coefficient and the correct invariant solution is proposed, but the scheme requires the same smoothness assumptions on $f$ and is numerically inferior to just simulating overdamped Langevin diffusion. We demonstrate our findings on both small-scale examples in which we can exactly verify the theoretical results, and on typical examples of larger scale in the context of Bayesian imaging inverse problems.
@@ 1 Introduction
$\mathrm{X}_{t}=\nabla\log\mu^{\ast}(x)\,\mathrm{d}t+\sqrt{2}\,\mathrm{d}W_{t},$ (1.2)
@@ 1.2 Contributions
  • We show that the primal marginal of the stationary solution in continuous time is not the target distribution. This can be interpreted as being a consequence of the dynamics not concentrating in the dual variable, although the diffusion coefficient is degenerate with the stochastic term only added to the primal variable.
@@ 2.1 Motivation for ULPDA
2.2 Continuous time persepctive
@@ 5 Numerical Examples
In the following, we demonstrate the behaviour and our new results on ULPDA on several numerical examples, both small-scale to demonstrate the convergence behaviour in detail and larger imaging inverse problems. We always simulate (2.5) in a time interval $[0,T]$ by running the discretized algorithm (2.4) with small step sizes $\tau,\sigma$ for $T/\tau$ steps. The step sizes are coupled via $\lambda\tau=\sigma$ and are chosen small enough so that $\tau\sigma L^{2}\leq 1$, i.e. the assumptions of Theorem 4.1 are satisfied and the continuous-time dynamics are approximated well.
@@ 5.2 TV Denoising from Gaussian Noise
For the potential $f\circ K$ which is dualized in the algorithm, we now consider the total variation (TV) functional, a typical model-based prior in imaging inverse problems. The (isotropic) TV functional is defined by
@@
where $\nabla_{h},\nabla_{v}$ denote horizontal and vertical finite differences in the pixelated image $x$. We consider a Gibbs prior on $x$ with density $\exp(-\alpha\operatorname{TV}(x))$ where $\alpha$ is a fixed regularization parameter (Note that this does not define a probability distribution since TV is invariant to constant translations and hence has a one-dimensional null space. With a log-likelihood that grows fast enough at infinity along this null space, the posterior is however well-defined.). The corresponding potential term can be written in the form $\alpha\operatorname{TV}(x)=\alpha\lVert Kx\rVert_{2,1}=f(Kx)$ where $K=(\nabla_{h},\nabla_{v})$ and $f=\alpha\lVert\cdot\rVert_{2,1}$ denotes a constant multiple of the nested $l_{2}$-$l_{1}$-norm (inner $l_{2}$ over vertical/horizontal component, outer $l_{1}$ over all pixels). We further assume that we make a noisy observation $\tilde{x}=x+\varepsilon$ of $x$, where the noise is distributed like $\varepsilon\sim\mathrm{N}(0,\sigma_{\varepsilon}^{2}I)$ for some observation variance $\sigma_{\varepsilon}^{2}$. The corresponding negative log-likelihood is
@@ 5.2.1 Toy Example in Two Dimensions
Using numerical integration, we are in this low-dimensional example able to compute an estimate of the normalization constant $Z=\int_{\mathcal{X}}\exp(-f(Kx)-g(x))\,\mathrm{d}x$. Using this, we can access an approximation of the posterior density $\mu^{\ast}(x)$. This allows us to compute approximations of the Wasserstein distances of current iterates to the target $\mathcal{W}_{2}(\hat{\mu}^{n},\mu^{\ast})$ without the need to rely on a second sampler. We evaluate the distance at 20 equidistantly spaced points in $[0,T]$ (i.e. for $\hat{\mu}^{kN},k=1,\dots,20$ where $N=T/20\tau$). The distances are computed using the Python optimal transport module [Flamary2021] and are displayed in Figure 4. As expected, the samples drawn by the Prox-Sub algorithm are closest to the target distribution (their bias in Wasserstein distance is bounded by a constant multiple of the step size $\tau$). For finite $\lambda$, the distributions $\hat{\mu}^{n}$ exhibit an additional strong bias to the target due to the dualization, which becomes smaller as $\lambda$ grows.
@@ 5.2.2 Image Denoising
We now sample from the posterior of an image denoising task. Due to the higher dimension, we can not compute accurate approximations of the Wasserstein distances to the target anymore. We therefore resort to illustrating the behaviour of ULPDA using moment estimates. A ground truth image normalized to intensities $[0,1]$ is distorted by noise with standard deviation $0.25$ (PSNR 12.0212.0212.02\,12.02dB in the noisy observation). We then draw $10^{6}$ samples with a single chain and compute the sample mean and pixel-wise variance after a burn-in phase of $31e$ 3 samples. The experiment is repeated three times for values $\lambda=1,10,100$, each with the same step size $\tau$. A fourth chain is computed using Prox-Sub with the same $\tau$. The sample mean (MMSE) estimates produced by the different chains converge quickly and are very similar for all chains (PSNR values between 23.8223.8223.82\,23.82dB and 23.8523.8523.85\,23.85dB). The pixel-wise variance, however, shows the same effect of overdispersion for smaller levels of $\lambda$. MMSE and pixel-wise variance estimates are reported in Figure 5.
@@
While the ‘primal’ samples $X^{n}$ are overdispersed for finite $\lambda$, the dual iterates $Y^{n}$ characteristically exhibit underdispersion (see Figure 6). This effect was consistent throughout all our numerical experiments (compare also Figures 2 and 3). A heuristic explanation is the fact that, since the maximization step in the dual variable solves the optimality condition $y\in\partial f(Kx)$ only incompletely, when the primal variable is driven towards the tail regions of the primal marginal distribution, the dual variable tends to remain around the mode of the dual marginal. The rate of overdispersion always decreased monotonically with growing parameters $\lambda$.

### v2

@@ Abstract
We analyze a recently proposed algorithm for the problem of sampling from probability distributions $\mu^{\ast}$ in $\mathbb{R}^{d}$ with a Lebesgue density of the form $\mu^{\ast}(x)\propto\exp(-f(Kx)-g(x))$, where $K$ is a linear operator and $f,g$ convex and non-smooth. The algorithm is a generalization of the primal-dual hybrid gradient (PDHG) convex optimization algorithm to a sampling scheme. We analyze the method’s continuous time limit, an SDE in the joint primal-dual variable. We give mild conditions under which the corresponding Fokker-Planck equation converges to a unique stationary state, which however does not concentrate in the dual variable and consequently does not have $\mu^{\ast}$ as its primal marginal. Under a smoothness assumption on $f$, we show that the scheme converges to the purely primal overdamped Langevin diffusion in the limit of small primal and large dual step sizes. We further prove that the target can never be the primal marginal of the invariant solution for any modification of the SDE with space-homogeneous diffusion coefficient. A correction with inhomogeneous diffusion coefficient and the correct invariant solution is proposed, but the scheme requires the same smoothness assumptions on $f$ and is numerically inferior to overdamped Langevin diffusion. We demonstrate our findings numerically first on small-scale examples in which we can exactly verify the theoretical results, and subsequently on typical examples of larger scale from Bayesian imaging inverse problems.
@@ 1 Introduction
$\mathrm{X}_{t}=\nabla\log\mu^{\ast}(X_{t})\,\mathrm{d}t+\sqrt{2}\,\mathrm{d}W_{t},$ (1.2)
@@ 1.2 Contributions
  • We show that the marginal with respect to the primal variable of the stationary solution in continuous time is not the target distribution. This can be interpreted as being a consequence of the dynamics not concentrating in the dual variable, although the diffusion coefficient is degenerate with the stochastic term only added to the primal variable.
@@ 2.1 Motivation for ULPDA
2.2 Continuous time perspective
@@ 5 Numerical Examples
In the following, we demonstrate the behaviour and our new results on ULPDA on several numerical examples, first small-scale ones to demonstrate the convergence behaviour in detail and subsequently larger imaging inverse problems. We always simulate (2.5) in a time interval $[0,T]$ by running the discretized algorithm (2.4) with small step sizes $\tau,\sigma$ for $T/\tau$ steps. The step sizes are coupled via $\lambda\tau=\sigma$ and are chosen small enough so that $\tau\sigma L^{2}\leq 1$, i.e. the assumptions of Theorem 4.1 are satisfied and the continuous-time dynamics are approximated well.
@@ 5.2 TV Denoising from Gaussian Noise
For the potential $f\circ K$ being is dualized in the algorithm, we now consider the total variation (TV) functional, a typical model-based prior in imaging inverse problems. The (isotropic) TV functional is defined by
@@
where $\nabla_{h},\nabla_{v}$ denote horizontal and vertical finite differences in the pixelated image $x$. We consider a Gibbs prior on $x$ with density $\exp(-\alpha\operatorname{TV}(x))$ where $\alpha$ is a fixed regularization parameter (Note that this does only defines a probability distribution on a subspace orthogonal to constant functions since TV has a one-dimensional null space. With a log-likelihood that grows fast enough at infinity along this null space, the posterior is however well-defined.). The corresponding potential term can be written in the form $\alpha\operatorname{TV}(x)=\alpha\lVert Kx\rVert_{2,1}=f(Kx)$ where $K=(\nabla_{h},\nabla_{v})$ and $f=\alpha\lVert\cdot\rVert_{2,1}$ denotes a constant multiple of the nested $l_{2}$-$l_{1}$-norm (inner $l_{2}$ over vertical/horizontal component, outer $l_{1}$ over all pixels). We further assume that we make a noisy observation $\tilde{x}=x+\varepsilon$ of $x$, where the noise is distributed like $\varepsilon\sim\mathrm{N}(0,\sigma_{\varepsilon}^{2}I)$ for some observation variance $\sigma_{\varepsilon}^{2}$. The corresponding negative log-likelihood is
@@ 5.2.1 Toy Example in Two Dimensions
Using numerical integration, in this low-dimensional example we are able to compute an estimate of the normalization constant $Z=\int_{\mathcal{X}}\exp(-f(Kx)-g(x))\,\mathrm{d}x$. Using this, we can access an approximation of the posterior density $\mu^{\ast}(x)$. This allows us to compute approximations of the Wasserstein distances of current iterates to the target $\mathcal{W}_{2}(\hat{\mu}^{n},\mu^{\ast})$ without the need to rely on a second sampler. We evaluate the distance at 20 equidistantly spaced points in $[0,T]$ (i.e. for $\hat{\mu}^{kN},k=1,\dots,20$ where $N=T/20\tau$). The distances are computed using the Python optimal transport module [Flamary2021] and are displayed in Figure 4. As expected, the samples drawn by the Prox-Sub algorithm are closest to the target distribution (their bias in Wasserstein distance is bounded by a constant multiple of the step size $\tau$). For finite $\lambda$, the distributions $\hat{\mu}^{n}$ exhibit an additional strong bias to the target due to the dualization, which becomes smaller as $\lambda$ grows.
@@ 5.2.2 Image Denoising
We now sample from the posterior of an image denoising task. Due to the higher dimension, we can not compute accurate approximations of the Wasserstein distances to the target anymore. We therefore resort to illustrating the behaviour of ULPDA using moment estimates. A ground truth image normalized to intensities $[0,1]$ is distorted by noise with standard deviation $0.25$ (PSNR 12.0212.0212.02\,12.02dB in the noisy observation). We then draw $10^{6}$ samples with a single chain and compute the sample mean and pixel-wise variance after a burn-in phase of $31e$ 3 samples. The experiment is repeated three times for values $\lambda=1,10,100$, each with the same step size $\tau$. A fourth chain is computed using Prox-Sub with the same $\tau$. MMSE and pixel-wise variance estimates are reported in Figure 5. The sample mean (MMSE) estimates produced by the different chains converge quickly and are very similar for all chains (PSNR values between 23.8223.8223.82\,23.82dB and 23.8523.8523.85\,23.85dB). The pixel-wise variance, however, shows the same effect of overdispersion for smaller levels of $\lambda$ as in the lower-dimensional example.
While the ‘primal’ samples $X^{n}$ are overdispersed for finite $\lambda$, the dual iterates $Y^{n}$ characteristically exhibit underdispersion (see Figure 6). This effect was consistent throughout all our numerical experiments (compare also Figures 2 and 3). A heuristic explanation for overdispersion in the primal and underdispersion in the dual variable lies in the fact that the step in the dual variable solves the optimality condition $y\in\partial f(Kx)$ only incompletely. On the particle level in continuous time, when the primal variable $X_{t}$ is driven towards the tail regions of its marginal distribution by the Brownian motion, the dual variable is not evolved entirely to the corresponding position $\partial f(KX_{t})$ in the tails of its marginal, but rather remains near the mode of its marginal. The effect decreased with growing parameters $\lambda$ in all our experiments.